Ask
your own question, for FREE!
Mathematics
19 Online
Heads I win, Tails you lose
Still Need Help?
Join the QuestionCove community and study together with friends!
John has 3 unique coins such that the probability of obtaining a head when the coin is flipped is \[\frac{1}{10},\frac{2}{10},\frac{ 3}{10}\], respectively. If he flips each of the 3 coins once, the probability that at least 1 heads appears is p. What is the value of \[1000p\]
If I understand this problem correctly, then they want you to use \[P(E)=\text{At least 1 heads in 3} \\ P(\overline{E})=\text{Maximum 0 heads in 3} \] Then \[\Large P(E)=1-P(\overline{E}) \]
|dw:1359988337180:dw|
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal